OpenStudy (anonymous):

the number of Math homework questions is seven more than one−half the number of History homework questions. There are 15 Math homework questions. Which of the following equations would you use to find the number of History homework questions? Note: h represents the number of questions on your History homework. 1. 15/2+7=h 2. 15+7/2=h 3. h+7/2=15 4. 1/2h+7=15 wish one is the correct answer

OpenStudy (anonymous):

Let the history questions be : h (as it is given).. Now, Maths questions = 7 + (1/2) of History Questions Maths questions are 15 and History questions are : h So, put these in the equation: \[15 = 7 + \frac{h}{2}\] Now can you get your answer from here??

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

my answer is 1/2 h + 7 =15 but i'm still not sure, so is it correct

OpenStudy (anonymous):

See, \[\frac{1}{2} h = \frac{h}{2}\] These are one and the same thing.. All right??

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

so the answer is h+7/2 = 15

OpenStudy (anonymous):

No you are wrong...